# Synthetic Lyrics Detection Across Languages and Genres

### v2

@@ Abstract
In recent years, the use of large language models (LLMs) to generate music content, particularly lyrics, has gained in popularity. These advances provide valuable tools for artists and enhance their creative processes, but they also raise concerns about copyright violations, consumer satisfaction, and content spamming. Previous research has explored content detection in various domains. However, no work has focused on the modality of lyrics in music. To address this gap, we curated a diverse dataset of real and synthetic lyrics from multiple languages, music genres, and artists. The generation pipeline was validated using both humans and automated methods. We conducted a comprehensive evaluation of existing synthetic text detection features on this novel data type. Additionally, we explored strategies to adjust the best feature for lyrics using unsupervised adaptation. Adhering to constraints of our application domain, we investigated cross-lingual generalization, data scalability, robustness to language combinations, and the impact of genre novelty in a few-shot detection scenario. Our findings show promising results within language families and similar genres, yet challenges persist with lyrics in languages that exhibit distinct semantic structures.
@@ 1 Introduction
Data, pre-processing scripts, code, and models will be publicly accessible under the Apache 2.0 license and in compliance with the content copyrights.
@@ 2 Related Work
As discussed above, previous research has explored content detection across various domains, yet no work has exclusively focused on the modality of lyrics in music. Moreover, prior benchmarks have primarily targeted English text and often lacked a rigorous validation of the synthetic text used in experiments, raising concerns about the findings’ reliability and generalization. These gaps highlight the need for a validated pipeline to generate and refine lyrics, the release of synthetic data that is realistic, musically diverse, and multilingual, and more targeted generalization experiments that explore various factors, including generative models, languages, and writing styles.
@@ 3.1 Human-Written Lyrics Dataset
For this work, we curated a multilingual dataset of 3,704 human-written lyrics targeting nine languages: English (EN), German (DE), Turkish (TR), French (FR), Portuguese (PT), Spanish (ES), Italian (IT), Arabic (AR), and Japanese (JA). The inclusion criterion was based on popularity, specifically from tracks listed in the most popular editorial playlists on an international music streaming platform222deezer.com as of June 2024. Also, we ensured that each track was released within the past year and a half to minimize the possibility that the models used in the detectors had prior exposure to this content. We evenly selected lyrics only from top-trending music genres per language, as determined by daily streaming statistics at extraction time. Appendix A shows the data distribution, and Appendix B the list of popular genres per language.
To allow a quality assessment of the generated lyrics by English-speaking humans from our organization, we decided to evenly and randomly pick a sub-sample from this dataset focused on the five most popular artists from the 2023 Billboard “Top Artists”333billboard.com/charts/year-end/top-artists, namely: Drake, Ed Sheeran, Post Malone, Taylor Swift, and The Weeknd. Though limited in scope, this dataset is a test bed of 625 human-written lyrics (for the distribution, see Appendix A) well-suited for assessing artistic style cloning capabilities of our LLM generation pipeline. We also use this controlled subset to identify the best detection features before running extensive experiments on robustness, scalability, and generalization.
@@ 3.2 Synthetic Lyrics Dataset
We opted for a constrained generation with a carefully designed prompt that was short and general, including some basic formatting instructions and three lyrics examples. The few-shot examples changed at each generation to diversify the output (Lu et al., 2022) but were conditioned on the same artist for the Billboard top artists data or the same language/genre pair for the multilingual data. To ensure the generated lyrics closely resembled real ones, the model was instructed to follow the same formatting guidelines as the real lyrics444docs.lyricfind.com. Appendix C shows the prompt template and Appendix D the hyperparameters used.
@@ 4.2 Detection Features
Perplexity (PPL) Beresneva (2016) measures the overall likelihood of the lyrics based on the exponential average of the negative log-likelihood. In principle, lower PPL suggests the lyrics are less likely to be human-written, although artistic writing could lead to higher PPL due to its unexpectedness.
@@ 5.3 Towards Evaluating Domain Adaptation
As shown in Figure 1, MNTP-based domain adaptation appears to outperform the base LLM2Vec model with no adaptation to the lyrics domain, with the gap seeming to increase with the size of the training dataset. The difference is particularly stark in some languages, such as Japanese, as shown in Appendix M. Our domain-adapted models are publicly available at HuggingFace.
@@ Methods
Appendix M Per-language Domain Adaptation Results

### v4

@@ Abstract
In recent years, the use of large language models (LLMs) to generate music content, particularly lyrics, has gained in popularity. These advances provide valuable tools for artists and enhance their creative processes, but they also raise concerns about copyright violations, consumer satisfaction, and content spamming. Previous research has explored content detection in various domains. However, no work has focused on the text modality, lyrics, in music. To address this gap, we curated a diverse dataset of real and synthetic lyrics from multiple languages, music genres, and artists. The generation pipeline was validated using both humans and automated methods. We performed a thorough evaluation of existing synthetic text detection approaches on lyrics, a previously unexplored data type. We also investigated methods to adapt the best-performing features to lyrics through unsupervised domain adaptation. Following both music and industrial constraints, we examined how well these approaches generalize across languages, scale with data availability, handle multilingual language content, and perform on novel genres in few-shot settings. Our findings show promising results that could inform policy decisions around AI-generated music and enhance transparency for users.
@@ 1 Introduction
Data, pre-processing scripts, code, and models will be publicly accessible on GitHub  [^2] https://github.com/deezer/synthetic_lyrics_detection under the Apache 2.0 license and in compliance with the content copyrights.
@@ 2 Related Work
As discussed above, previous research has explored content detection across various domains, yet no work has exclusively focused on the text modality, lyrics, in music. Moreover, prior benchmarks have primarily targeted English text and often lacked a rigorous validation of the synthetic text used in experiments, raising concerns about the findings’ reliability and generalization. These gaps highlight the need for a validated pipeline to generate and refine lyrics, the release of synthetic data that is realistic, musically diverse, and multilingual, and more targeted generalization experiments that explore various factors, including generative models, languages, and writing styles.
@@ 3.1 Human-Written Lyrics Dataset
For this work, we curated a multilingual dataset of 3,704 human-written lyrics targeting nine languages: English (EN), German (DE), Turkish (TR), French (FR), Portuguese (PT), Spanish (ES), Italian (IT), Arabic (AR), and Japanese (JA). The inclusion criterion was based on popularity, specifically from tracks listed in the most popular editorial playlists on an international music streaming platform333deezer.com as of June 2024. Also, we ensured that each track was released within the past year and a half to minimize the possibility that the models used in the detectors had prior exposure to this content. We evenly selected lyrics only from top-trending music genres per language, as determined by daily streaming statistics at extraction time. Appendix A shows the data distribution, and Appendix B the list of popular genres per language.
To allow a quality assessment of the generated lyrics by English-speaking humans from our organization, we decided to evenly and randomly pick a sub-sample from this dataset focused on the five most popular artists from the 2023 Billboard “Top Artists”444billboard.com/charts/year-end/top-artists, namely: Drake, Ed Sheeran, Post Malone, Taylor Swift, and The Weeknd. Though limited in scope, this dataset is a test bed of 625 human-written lyrics (for the distribution, see Appendix A) well-suited for assessing artistic style cloning capabilities of our LLM generation pipeline. We also use this controlled subset to identify the best detection features before running extensive experiments on robustness, scalability, and generalization.
@@ 3.2 Synthetic Lyrics Dataset
We opted for a constrained generation with a carefully designed prompt that was short and general, including some basic formatting instructions and three lyrics examples. The few-shot examples changed at each generation to diversify the output (Lu et al., 2022) but were conditioned on the same artist for the Billboard top artists data or the same language/genre pair for the multilingual data. To ensure the generated lyrics closely resembled real ones, the model was instructed to follow the same formatting guidelines as the real lyrics555docs.lyricfind.com. Appendix C shows the prompt template and Appendix D the hyperparameters used.
@@ 4.2 Detection Features
Perplexity (PPL) Beresneva (2016) measures the overall likelihood of the lyrics based on the exponential average of the negative log-likelihood. In principle, lower PPL suggests the lyrics are less likely to be human-written as artistic writing could lead to higher PPL due to its unexpectedness.
@@ 5.3 Towards Evaluating Domain Adaptation
As shown in Figure 1, MNTP-based domain adaptation appears to outperform the base LLM2Vec model with no adaptation to the lyrics domain, with the gap seeming to increase with the size of the training dataset. The difference is particularly stark in some languages, such as Japanese, as shown in Appendix P.
@@ Methods
Appendix M Effect of $k$ in kNN
We have chosen the best K experimentally on a smaller validation set from the Billboard, English only data. In Table 16, we show results on the multilingual test corpus when using the LLM2Vec embeddings (we could notice a similar trend for the other detection features). Similar to the behaviour on the English-only dataset, increasing the K higher than 3 does not increase the scores much.
Langs k=1 k=3 k=5 k=10 k=20
EN 90.97 89.55 89.75 89.41 72.49
DE 97.46 97.61 98.07 98.14 98.17
TR 82.54 82.76 82.76 82.76 82.54
FR 96.84 97.71 98.14 98.14 98.15
PT 89.28 89.46 89.22 90.76 90.72
ES 94.11 92.33 92.22 92.00 92.11
IT 80.53 83.09 82.79 82.58 80.91
AR 92.01 92.03 92.62 92.81 91.43
JA 70.43 70.85 71.23 69.34 70.47
Avg. 88.24 88.38 88.53 88.44 86.33
Table 16: Results on the multilingual dataset with LLM2Vec + Llama3 8B when varying $k$ in kNN.
Appendix N Results with AUROC
Language LLM2Vec LUAR Entropy PPL
EN 96.5 100.0 99.1 63.3
DE 98.0 99.5 97.4 61.0
TR 92.9 92.9 68.4 58.7
FR 99.4 99.0 98.2 66.1
PT 97.1 99.6 99.6 60.2
ES 95.1 99.6 96.9 55.4
IT 90.4 100.0 95.2 60.8
AR 93.7 95.9 68.9 62.1
JA 80.7 94.1 87.4 59.4
Avg. 93.7 97.8 90.1 60.8
Table 17: Results on the multilingual dataset with AUROC using four different classifiers.
The AUROC analysis reveals distinct patterns across detection methods and languages. LUAR demonstrates superior performance (97.8% average), particularly excelling in Indo-European languages with perfect or near-perfect scores. While LLM2Vec (93.7% average) and the Entropy-based classifier (90.1%) perform well on Indo-European languages, they struggle significantly with morphologically rich languages like Turkish and Arabic (around 68% for Entropy) and different writing systems like Japanese (80.7% for LLM2Vec). The Perplexity-based approach’s consistent underperformance (60.8% average) across all languages suggests fundamental limitations in using raw probability scores for detection.
Appendix O Results with Majority Voting Classifier
LLM2Vec Max NLL UAR Entropy Maj.
EN 90.6 59.3 100.0 96.6 100.0
DE 97.4 56.7 99.2 97.2 99.0
TR 82.7 56.5 82.4 65.0 82.5
FR 97.7 62.1 97.6 97.8 97.9
PT 89.2 54.8 99.3 99.1 99.4
ES 92.3 54.7 99.0 95.0 97.3
IT 83.0 63.3 100.0 95.9 98.0
AR 92.1 58.9 93.6 65.8 93.8
JA 71.5 55.3 74.6 86.2 78.7
Avg. 88.5 58.0 94.0 88.7 94.1
Table 18: Per feature performances over all languages for the baseline scenario with a majority voting classifier, combining votes from the 4 best-performing classifiers, which are also shown for clarity.
The majority voting approach (Maj.) achieves the highest average performance at 94.1%, showing only marginal improvement over UAR at 94.0%. This minimal gain suggests that combining multiple classifiers through majority voting does not provide substantial benefits over the best individual classifier (UAR). The similar performance between majority voting and UAR also suggests that the different detection methods might be capturing similar features or making correlated errors, limiting the potential benefits of ensemble approaches.
Appendix P Per-language Domain Adaptation Results